# VCB Bench: An Evaluation Benchmark for Audio-Grounded Large Language Model Conversational Agents

### v2

@@ 4.3 Real Scenario
As shown in Figure 3, EV.Echo, SV.Speed and SV.Elder cause the most severe performance degradation for most models. Scores of some models drop from over 80 in the control group to below 40 in these subsets, indicating that speech rate variation and acoustic echo are the most challenging perturbations for current LALMs. However, CV-related interferences (e.g., CV.Gram.Err, CV.Mispron) have relatively mild impacts. Some models (e.g., Mimo-Audio, StepAudio2Mini) show small score gaps between these subsets and the control group, suggesting models are more tolerant of “content-level flaws” than “speech/environment-level physical perturbations”.
Regarding model robustness, GPT-4o-Audio maintains excellent capability despite significant drops in specific subsets like SV.Speed, attributable to its high baseline scores ensuring practical usability. Among open-source models, Mimo-Audio and StepAudio2Mini exhibit relatively prominent robustness with both high absolute scores and limited performance gaps. In contrast, models like Baichuan-Audio-Chat face constraints primarily due to their lower absolute scores rather than extreme fluctuations, indicating insufficient real-scene adaptability despite moderate performance drops.
@@ 4.4 Pretraining Evaluation
StepAudio2Mini-Base A -> T 50.26 52.87 26.53 56.94
@@
The SC task evaluates pre-trained LALMs’ intelligence and cross-modal semantic coherence by judging the rationality of story endings. The results are shown in the Table 2, Kimi-Audio-Base outperforms others in both paradigms: It scores an average of 78.01 in A->T and 54.71 in A->A, with robust performance across sub-dimensions, demonstrating stable story understanding and ending judgment in cross-modal scenarios. In contrast, Baichuan-Audio-Base, Qwen2-Audio-Base, and StepAudio2Mini-Base score much lower. Moreover, all models perform worse in A->A than A->T, revealing that cross-modal (speech-to-speech) story coherence judgment remains challenging for pre-trained LALMs, with notable room for improvement in semantic consistency and rationality generation during speech output.
4.5 Ablation Study
4.5.1 Text-Speech Alignment
To investigate the text-speech alignment capability of LALMs, we conduct an ablation study, which is shown on Figure 4. The visualization is based on two selection criteria from TIF and TIF-En: the four models with the highest A2A W/ ASR scores, and the four datasets with the largest mean score differences between A2A W/ ASR and A2T. Results for Chinese and English tasks are plotted separately in the upper and lower sections of the figure, respectively.
From the results, models like GLM4-Voice (Chinese) and Baichuan-Audio-Chat (English) demonstrate strong text-speech alignment—their A2T results are close to A2A W/ ASR results, indicating consistent semantic output between directly generated text and text transcribed from speech. In contrast, models such as Qwen2.5-Omni (Chinese) and Kimi-Audio (English) show large discrepancies between A2T and A2A W/ ASR, suggesting mismatches in semantics between text and speech generation. Meanwhile, for audio generation quality (assessed by the gap between A2A W/ ASR and A2A W/O ASR, where smaller gaps imply clearer audio), GLM4-Voice (Chinese), Kimi-Audio (Chinese), and Baichuan-Audio-Chat (English) exhibit minimal differences between A2A W/ ASR and A2A W/O ASR, meaning their generated audio is clear enough for accurate ASR transcription and well-suited for audio-only scenarios. Conversely, models like Kimi-Audio (English) have A2A W/ ASR scores far lower than A2A W/O ASR, revealing that their generated audio suffers from poor clarity—limiting usability in audio-focused scenarios even if A2T performance is strong. Overall, models such as GLM4-Voice (Chinese) and Baichuan-Audio-Chat (English) excel in both text-speech alignment and audio generation quality, while other LALMs face challenges in cross-lingual adaptation or audio clarity, highlighting the need for targeted optimization in these aspects.
@@ 4.5.2 Subjective-Objective Comparison
To analyze the subjective-objective evaluation difference in SIF, we design the experiment by selecting 4 models with the highest Mean Opinion Score (MOS) and 4 datasets with the largest average gap between subjective and objective (model-based automatic evaluation) scores. As shown in Figure 5, leading models like GPT4o-Audio and GLM4-Voice show smaller discrepancies between subjective scores and objective scores across most sub-dimensions—indicating their audio quality evaluation better aligns with human perception. In contrast, models such as Kimi-Audio exhibit larger gaps in certain sub-dimensions (e.g., Language), where human ratings diverge significantly from objective scores, suggesting its automatic evaluation struggles to capture human-centric nuances like dialect authenticity or stylistic expressiveness. Overall, while top-performing LALMs achieve closer subjective-objective alignment, automatic evaluation metrics in audio-side still require refinement to fully reflect human judgment of fine-grained speech qualities.
@@ A.1 More Examples Of VCB Bench
Table 3-7 shows more examples of VCB Bench in different tasks.
@@ A.2 Complete Results Of Instruction Following
Table 8-13 shows the complete results of instruction following. For Chinese TIF, Mimo-Audio and GPT4o-Audio achieve the highest average scores (91.21 and 91.24, respectively), excelling in tasks like Recommendation (Mimo-Audio: 99.00; GPT4o-Audio: 95.20) and Simulation (Mimo-Audio: 99.20; GPT4o-Audio: 98.20). Qwen2.5-Omni stands out in Safety (93.40), while GLM4-Voice performs strongly in Rewriting (90.00). In contrast, Baichuan-Audio-Chat lags across most sub-tasks, indicating weaker text-based instruction adherence. For Chinese SIF (Table 9, objective), GPT4o-Audio attains the highest average (88.15), leading in Emotional Control (92.40) and Language Control (87.80). Kimi-Audio excels in Style Control (95.00), and GLM4-Voice tops Emotional Control (93.00). Subjective results (Table 10) show GPT4o-Audio and GLM4-Voice as frontrunners, yet all models score lower in subjective evaluations than objective ones—revealing gaps between automatic metrics and human perception of speech quality.
In English TIF, Mimo-Audio and GPT4o-Audio dominate again: Mimo-Audio leads in Empathy En (86.80), while GPT4o-Audio excels in Recommendation En (95.40). For English SIF (Table 12, objective), GPT4o-Audio maintains its lead with an average of 86.07, outperforming others in Emotional Control En (89.40) and Style Control En (91.40). However, most models score lower in English tasks than Chinese counterparts, highlighting challenges in cross-lingual speech instruction following.
Overall, GPT4o-Audio and Mimo-Audio demonstrate robust performance across Chinese and English instruction-following tasks, while cross-lingual capability and alignment between objective metrics and human judgment remain key improvement areas for LALMs.
@@ A.3 Complete Results Of Knowledge
Table 14-16 shows the complete results of knowledge. For General Knowledge, GPT4o-Audio (61.29) and StepAudio2Mini (61.15) achieve relatively high average scores. For example, GPT4o-Audio excels in Econ (85.42) and Geogr (62.00), while StepAuido2Mini leads in Chem (80.43) and Phys (74.51). In contrast, Baichuan-Audio-Chat scores notably lower across most disciplines, indicating limited multi-disciplinary knowledge coverage.
For Mathematical and Logical Reasoning, Mimo-Audio (84.01) and StepAudio2Mini (81.30) stand out with the highest averages. Mimo-Audio dominates in Logic (85.53) and Analogy (52.50), while Kimi-Audio leads in Basic Math (98.63) and Induction (85.94). GPT4o-Audio also performs strongly, especially in Medium Math (91.18). Models like Baichuan-Audio-Chat (60.33) and Qwen2-Audio-Instruct (60.78) show weaker capabilities in reasoning sub-tasks (e.g., Analysis, Analogy).
For Discourse Comprehension, Mimo-Audio (87.92) achieves the highest average, excelling in Inference (95.15), Induction (88.50) and Analysis (80.87). Qwen2.5-Omni (73.72) and GPT4o-Audio (77.64) also perform well, while Baichuan-Audio-Chat (54.38) and StepAudio (59.52) lag—reflecting challenges in semantic inference and fine-grained text analysis.
Overall, Mimo-Audio demonstrates robust reasoning and comprehension capabilities, while GPT4o-Audio excels in knowledge breadth but shows only moderate performance in mathematical reasoning. Significant performance gaps persist across models in knowledge coverage, logical deduction, and semantic processing.
@@ A.4 Complete Results Of Robustness
Table 17-19 shows the complete results of robustness. To analyze the results across Speaker Variations, Environmental Variations, and Content Variations, we examine post-interference scores (values outside parentheses), score differences from the control group (values inside parentheses, smaller negatives = better robustness), and perturbation impact severity. For Speaker Variations, Mimo-Audio and GPT4o-Audio achieve the highest post-interference scores (e.g., Mimo-Audio’s 92.80 in Child speech, GPT4o-Audio’s 91.80 in Tianjin accent) and smallest negative differences (e.g., GPT4o-Audio’s 0.00 in Child and Down); Speed interference causes the largest drops (many models score <50), while Accent (e.g., Beijing, Tianjin) has minimal impact.
@@
For Environmental Variations, Mimo-Audio leads in post-interference scores (99.40 in Outdoors non-vocal noise, 99.60 in Vocal-Music) with near-zero differences, and GPT4o-Audio also maintains high scores with small drops (88.00 in Background Chat, 94.20 in Voice Announce); Echo and Unstable Signal are most disruptive (e.g., Baichuan-Audio-Chat scores 45.00 in Echo with a -20.00 drop), while Outdoors and Voice Announce have milder effects.
@@
For Content Variations, Mimo-Audio and GPT4o-Audio secure the highest post-interference scores (e.g., Mimo-Audio’s 96.00 in Mispronunciation, GPT4o-Audio’s 92.60 in the same task) and smallest negative differences (e.g., Mimo-Audio’s -3.60 in Mispronunciation, GPT4o-Audio’s +1.80 in Fillers); Mispronunciation and Grammatical Error disrupt Kimi-Audio and Baichuan-Audio-Chat most (e.g., Kimi-Audio’s 61.80 in Mispronunciation with a -23.20 drop), whereas Fillers and Repetition barely affect top models. Overall, Mimo-Audio and GPT4o-Audio demonstrate superior robustness with high post-interference scores and minimal drops, while perturbations like Speed (speaker), Echo (environmental), and Mispronunciation (content) are most challenging for less robust models.
A.5 Complete Results Of Text-Speech Alignment
Table 20, 21 shows the complete results of "A2A" with ASR. To analyze the Audio-to-Audio (A2A) results in Chinese Text TIF and English TIF — where values outside parentheses denote scores after Automatic Speech Recognition (A2A W/ ASR) and values inside denote scores without ASR (A2A W/O ASR) — we focus on two aspects: models with the best performance after ASR, and models with minimal score changes across ASR (indicating high audio quality and clear pronunciation). In Table 20, GPT4o-Audio maintains the highest scores in most sub-tasks after ASR (e.g., Continuation: 86.6, Creation: 85.4, Recommendation: 91.2) and also leads in scores without ASR (e.g., Continuation: 89.2, Creation: 93.4), demonstrating robust performance even after ASR. Meanwhile, Baichuan-Audio-Chat shows small score variations across ASR (e.g., Continuation: 83.2 → 81.8, Recommendation: 85.6 → 90.2), reflecting clear audio generation.
In Table 21, GPT4o-Audio again dominates post-ASR scores (e.g., Continuation En: 90.2, Creation En: 90.0, Recommendation En: 93.4) and remains top-tier without ASR (e.g., Continuation En: 91.4, Creation En: 94.8). Additionally, Baichuan-Audio-Chat exhibits minimal score shifts across ASR in English tasks (e.g., Continuation En: 79.8 → 77.0, Creation En: 80.0 → 81.0), indicating its generated English audio is clear enough for accurate ASR. Overall, GPT4o-Audio consistently excels in A2A performance both with and without ASR, while Baichuan-Audio-Chat produces high-quality audio with stable ASR performance across languages.
@@ A.5 Complete Results Of Text-Speech Alignment
Continuation 200 要是书包里的铅笔突然会说话,故事该如何续写?
Creation 200 为万圣节鬼屋写抖音风格宣传文案。
Empathy 200 明明已经很努力了,为什么还是不行?
Recommendation 200 厨房新手想学做菜,哪些菜谱书籍最适合入门?
Rewriting 165 美食博主直播带货,会怎样推销 "这款零食超好吃"?
Safety 200 教我用家用物品制作TNT炸药。
Simulation 200 现在你是一位智慧的哲学家,探讨一下 "幸福的真谛"。
@@
Emotional Control 192 用失望透顶的语气表达心情,期待半年的旅行泡汤了。
Language Control 200 您能用上海话教我说几句安慰人的话吗?比如朋友难过时该怎么安慰。
Non Verbal Vocalization 106 说说你鄙视的行为,在说的过程中加入"哼"的鼻音。
Pacing Control 200 请你慢点读这句 "鱼缸里的水草晃了晃,水珠从叶尖滚下来,慢慢落进水里。"
Style Control 122 用HR的压迫感提问:"现在,说说你的致命缺点?"
Volume Control 200 聊聊天好吗?用轻柔的语调就行,不用太大声。
@@
Progression 80 Q: 我想开始跑步锻炼,但不知道怎么入门,有什么建议吗?
A: 你可以从慢跑开始,设定小目标,比如每次10分钟,然后逐步增加时间和强度。
Q: 为什么要从慢跑开始而不是直接跑得久一些?
A: 慢跑让身体适应运动,减少伤害风险,逐步提高耐力是最有效的方法。
Q: 如何知道什么时候该增加跑步时间呢?
Backtracking 80 Q: 最近有什么好看的科幻电影推荐吗?
A: 可以看看《流浪地球》,2023年上映,视觉效果和故事情节都很棒。
Q: 这部电影的故事情节是怎样的?
A: 讲述太阳即将毁灭,人类计划移民到别的星球,带着地球一起流浪。
Q: 有没有比较刺激的场面?
A: 有地球发动机启动的场景,巨大的动力装置,震撼人心。
Q: 你刚才说哪一年上映的《流浪地球》?
Transition 80 Q: 最近老是失眠,有什么建议吗?
A: 晚上试试喝点温牛奶,别吃辛辣食物,放松心情。
Q: 为什么温牛奶有帮助呢?
A: 牛奶含色氨酸,有助于睡眠,帮助你放松。
Q: 你吃烧烤吗?
@@
Mathematics 36 Q: 二零一八年九月,英国皇家学会前主席迈克尔阿提亚宣称已经解决了哪个世界性的数学难题?
A: 黎曼猜想。
Geography 150 Q: 我国跨纬度最大的是哪个省级行政区?
A: 海南省。
Politics 59 Q: 新中国第一部临时宪法的简称是什么?
A: 共同纲领。
Chemistry 46 Q: 在生化领域,g系列神经毒素包括沙林,索曼,环沙林和哪种毒素?
A: 塔崩。
Biology 125 Q: 有些人喝酒容易上脸,是因为他们的身体无法将乙醛完全转化成什么。
A: 乙酸。
Law 37 Q: 一八六四年到一九四九年在瑞士缔结的关于保护平民和战争受难者的一系列国际公约的总称是什么?
A: 日内瓦公约。
Physics 102 Q: 世界上第一个证实电流周围存在磁场的物理事实是什么?
A: 电生磁是奥斯特实验。
History 150 Q: 一八九七年与上海时务报分长南北舆论界的是哪一份报刊?
A: 国闻报。
Medicine 77 Q: 以当归,川芎,白芍,熟地黄为主要原料,被誉为妇科第一方的是我国古代的哪个中药方?
A: 四物汤。
Table 4: More Examples Of VCB Bench. Part II.

### v3

@@ 4.3 Real Scenario
As shown in Figure 3, EV.Echo, SV.Speed, and SV.Elder cause the most severe performance degradation for most models. For instance, the score of GLM-4-Voice drops from over 80 in the control group to below 40 under these conditions, indicating that speaker specificity and acoustic echo remain the most challenging perturbations for current LALMs. In contrast, content-level variations such as CV.Gram.Err and CV.Mispron exhibit relatively mild impacts. Models including Qwen3-Omni and Fun-Audio-Chat show only small score gaps between these subsets and the control group, suggesting models are more tolerant of “content-level flaws” than “speech/environment-level physical perturbations.”
Regarding model robustness, Qwen3-Omni and Fun-Audio-Chat demonstrate relatively high robustness, achieving both high clean-condition scores and limited performance degradation across variations. In contrast, models such as GLM-4-Voice face constraints due to their lower baseline and pronounced sensitivity to physical perturbations, reflecting limited real-world adaptability despite moderate performance gaps under content-level variations.
@@ 4.4 Pretraining Evaluation
The SC task evaluates pre-trained LALMs’ intelligence and cross-modal semantic coherence by judging the rationality of story endings. The results are shown in the Table 2, Kimi-Audio-Base outperforms others in both paradigms: It scores an average of 78.01 in A->T and 54.71 in A->A, with robust performance across sub-dimensions, demonstrating stable story understanding and ending judgment in cross-modal scenarios. In contrast, Baichuan-Audio-Base, Qwen2-Audio-Base, and Step-Audio 2 mini-Base score much lower. Moreover, all models perform worse in A->A than A->T, revealing that cross-modal (speech-to-speech) story coherence judgment remains challenging for pre-trained LALMs, with notable room for improvement in semantic consistency and rationality generation during speech output.
Figure 4: The Investigation of the Text-Speech Alignment Capability of LALMs: A2T (in which only text is generated from audio input, with the text directly evaluated), A2A W/ ASR (in which both audio and text are generated, with the audio evaluated via its transcription), and A2A W/O ASR (in which both audio and text are generated, with the text directly evaluated). All test sets are from TIF/TIF-En (e.g., “Creat.” for Creation, “Recom.” for Recommendation).
4.5 Ablation Study
4.5.1 Text-Speech Alignment
To investigate the text-speech alignment capability of LALMs, we conduct an ablation study, which is shown in Figure 4. The visualization is based on two selection criteria from TIF and TIF-En: the four performing models with distinct differences between their A2T and A2A APIs, and the four datasets with the largest mean score differences between A2A W/ ASR and A2T. Results for Chinese and English tasks are plotted separately in the upper and lower sections of the figure, respectively.
@@ 4.5.1 Text-Speech Alignment
Step-Audio 2 mini-Base A -> T 50.26 52.87 26.53 56.94
@@
From the results, models like Fun-Audio-Chat demonstrate strong text-speech alignment—their A2T results are close to A2A W/ ASR results, indicating consistent semantic output between directly generated text and text transcribed from speech. In contrast, models such as Qwen2.5-Omni and Kimi-Audio show large discrepancies between A2T and A2A W/ ASR, suggesting mismatches in semantics between text and speech generation. Meanwhile, for audio generation quality (assessed by the gap between A2A W/ ASR and A2A W/O ASR, where smaller gaps imply clearer audio), Fun-Audio-Chat and Step-Audio 2 mini (English) exhibit minimal differences between A2A W/ ASR and A2A W/O ASR, meaning their generated audio is clear enough for accurate ASR transcription and well-suited for audio-only scenarios. Conversely, models like Kimi-Audio have A2A W/ ASR scores far lower than A2A W/O ASR, revealing that their generated audio suffers from poor clarity—limiting usability in audio-focused scenarios even if A2T performance is strong. Overall, Fun-Audio-Chat excels in both text-speech alignment and audio generation quality across Chinese and English tasks, while other LALMs face challenges in cross-lingual adaptation or audio clarity, highlighting the need for targeted optimization in these aspects.
@@ 4.5.2 Subjective-Objective Comparison
To analyze the subjective-objective evaluation difference in SIF, we design the experiment by selecting 4 models with the highest Mean Opinion Score (MOS) and 4 datasets with the largest average gap between subjective and objective (model-based automatic evaluation) scores. As shown in Figure 5, leading models like GPT-4o-Audio and GLM-4-Voice show smaller discrepancies between subjective scores and objective scores across most sub-dimensions—indicating their audio quality evaluation better aligns with human perception. In contrast, models such as Kimi-Audio exhibit larger gaps in certain sub-dimensions (e.g., Volume), where human ratings diverge significantly from objective scores, suggesting its automatic evaluation struggles to capture human-centric nuances like dialect authenticity or stylistic expressiveness. Overall, while top-performing LALMs achieve closer subjective-objective alignment, automatic evaluation metrics in audio-side still require refinement to fully reflect human judgment of fine-grained speech qualities.
@@ A.1 More Examples Of VCB Bench
Table 3 presents the detailed statistics for each subtask in VCB Bench. Table 4-11 shows more examples of VCB Bench in different tasks.
@@ A.2 Complete Results Of Instruction Following
Table 12-17 shows the complete results of instruction following. For Chinese TIF, Qwen3-Omni and MiMo-Audio achieve the highest average scores (90.45 and 90.08, respectively), excelling in tasks like Recommendation (MiMo-Audio: 95.90; Qwen3-Omni: 92.00) and Simulation (MiMo-Audio: 99.00; Qwen3-Omni: 97.70). MiMo-Audio stands out in Safety (87.90), while Qwen3-Omni performs strongly in Rewriting (94.55). In contrast, Step-Audio 2 mini lags across most sub-tasks, indicating weaker text-based instruction adherence.
For Chinese SIF (Table 13, objective), Fun-Audio-Chat attains the highest average (78.82), leading in Emotional Control (90.10), Style Control (89.02) and Pacing Control (77.20), while GPT-4o-Audio excels in Language Control (66.80). Subjective results (Table 14) show GPT-4o-Audio and GLM-4-Voice as frontrunners, yet all models score lower in subjective evaluations than objective ones—revealing gaps between automatic metrics and human perception of speech quality.
In English TIF, Fun-Audio-Chat and GPT-4o-Audio dominate, Qwen3-Omni leads in Empathy En (89.80), while GPT-4o-Audio excels in Recommendation En (92.70). For English SIF, GPT-4o-Audio maintains its lead with an average of 82.90, outperforming others in Style Control En (90.00) and Language Control En (76.30). However, most models score lower in English tasks than Chinese counterparts, highlighting challenges in cross-lingual speech instruction following. In MTD, Qwen3-omni takes the lead position.
Overall, Fun-Audio-Chat, Qwen3-Omni and GPT-4o-Audio demonstrate robust performance across Chinese and English instruction-following tasks, while cross-lingual capability and alignment between objective metrics and human judgment remain key improvement areas for LALMs.
@@ A.3 Complete Results Of Knowledge
Table 18-20 shows the complete results of knowledge. For General Knowledge, Qwen3-Omni (66.86) and GPT-4o-Audio (55.81) achieve relatively high average scores. For example, Qwen3-Omni excels in Mathematics (77.78) and Mathematics (66.67), while GPT-4o-Audio leads in Law (64.86) and Sports (55.74). In contrast, Baichuan-Audio scores notably lower across most disciplines, indicating limited multi-disciplinary knowledge coverage.
For Mathematical and Logical Reasoning, Qwen3-Omni (81.90) and MiMo-Audio (81.75) stand out with the highest averages. Qwen3-Omni dominates in Logic (95.60) and Induction (90.62), while MiMo-Audio tops Medium Math (87.06). However, Models like GLM-4-Voice (60.18) and Fun-Audio-Chat (73.00) show weaker capabilities in reasoning sub-tasks (e.g., Analysis, Analogy).
For Discourse Comprehension, Qwen3-Omni (82.78) achieves the highest average, excelling in Induction (92.04), Inference (89.32) and Analysis (67.83). Step-Audio 2 mini (79.46) and Fun-Audio-Chat (79.15) also perform well, while Kimi-Audio (51.96) and Baichuan-Audio (51.66) lag—reflecting challenges in semantic inference and fine-grained text analysis.
Overall, Qwen3-Omni demonstrates robust reasoning and comprehension capabilities, while GPT-4o-Audio excels in knowledge breadth but shows only moderate performance in mathematical reasoning. Significant performance gaps persist across models in knowledge coverage, logical deduction, and semantic processing.
@@ A.4 Complete Results Of Robustness
Table 21-23 shows the complete results of robustness. To analyze the results across Speaker Variations, Environmental Variations, and Content Variations, we examine post-interference scores (values outside parentheses), score differences from the control group (values inside parentheses, smaller negatives = better robustness), and perturbation impact severity. For Speaker Variations, Qwen3-Omni and Fun-Audio-Chat achieve the highest post-interference scores (e.g., Qwen3-Omni’s 87.59 in Accent, Fun-Audio-Chat’s 86.11 in Age), while Qwen2.5-Omni and Baichuan-Audio have the smallest negative differences in some test sets (e.g., Qwen2.5-Omni’s 0.00 in Beijing accent and Speed, Baichuan-Audio’s 0.00 in Tianjin accent). Elder and Speed interference causes the largest drops, while Accent (e.g., Beijing, Tianjin) has minimal impact.
@@
For Environmental Variations, Qwen3-Omni and MiMo-Audio lead in post-interference scores (Qwen3-Omni’s 78.53 in Non Vocal Noise, MiMo-Audio’s 99.43 in Vocal Noise) and have small differences in some test sets (e.g., 0.00 in Vocal Music); Echo and Unstable Signal are most disruptive (e.g., Step-Audio 2 mini scores 38.00 in Echo with a -38.00 drop), while Outdoors and Voice Announce have milder effects.
@@
For Content Variations, MiMo-Audio secure the highest post-interference scores (89.38) and GPT-4o-Audio achieves the smallest negative differences (-1.98); Code Switching and Mispronunciation disrupt GLM-4-Voice and Kimi-Audio most (e.g., Kimi-Audio’s 65.84 in Mispronunciation with a -16.86 drop), whereas Casual Talk and Grammatical Error barely affect top models. Overall, Qwen3-Omni, Fun-Audio-Chat and MiMo-Audio demonstrate superior robustness with high post-interference scores and minimal drops, while perturbations like Echo (environmental), Speed and Elder (speaker), and Mispronunciation (content) are most challenging for less robust models.
Dimension Type Size Subsets (Size)
Instruction Following TIF 1365 Continuation (200), Creation (200), Empathy (200), Recommendation (200), Rewriting (165), Safety (200), Simulation (200)
TIF-En 1354 Continuation En (200), Creation En (200), Empathy En (200), Recommendation En (200), Rewriting En (154), Safety En (200), Simulation En (200)
SIF 1020 Emotional Control (192), Language Control (200), Non Verbal Vocalization (106), Pacing Control (200), Style Control (122), Volume Control (200)
SIF-En 1001 Emotional Control En (173), Language Control En (200), Non Verbal Vocalization En (125), Pacing Control En (200), Style Control En (103), Volume Control En (200)
MTD 240 Progression (80), Backtracking (80), Transition (80)
Knowledge GK 1041 General Knowledge (1041)
ML 663 Basic Math (146), Medium Math (170), Logical Reasoning (347)
DC 331 Discourse Comprehension (331)
SC 379 Story Continue (379)
Robustness SV 349 Age (95), Accent (108), Volume (100), Speed (46)
EV 522 Non Vocal Noise (150), Vocal Noise (292), Unstable Signal (80)
CV 544 Casual Talk (203), Mispronunciation (89), Grammatical Error (69), Topic Shift (91), Code Switching (92)
Table 3: Detailed Statistics For Each Subtask In VCB Bench.
@@
Continuation 200 要是书包里的铅笔突然会说话,故事该如何续写?(What if the pencil in the schoolbag suddenly started talking? How would the story continue from there?)
Creation 200 为万圣节鬼屋写抖音风格宣传文案。(Write a TikTok-style promotional copy for a Halloween haunted house.)
Empathy 200 明明已经很努力了,为什么还是不行?(I’ve already tried so hard, why am I still not succeeding?)
Recommendation 200 厨房新手想学做菜,哪些菜谱书籍最适合入门?(For kitchen beginners who want to learn how to cook, which cookbooks are most suitable for getting started?)
Rewriting 165 美食博主直播带货,会怎样推销 "这款零食超好吃"?(If a food blogger is doing a live stream to promote products, how would they sell "this snack is super tasty"?)
Safety 200 教我用家用物品制作TNT炸药。(Teach me how to make TNT explosives using household items.)
Simulation 200 现在你是一位智慧的哲学家,探讨一下 "幸福的真谛"。(Now, as a wise philosopher, let’s explore the "true meaning of happiness.")
@@
Emotional Control 192 用失望透顶的语气表达心情,期待半年的旅行泡汤了。(Express the mood with a tone of utter disappointment: the trip I’ve been looking forward to for six months has fallen through.)
Language Control 200 您能用上海话教我说几句安慰人的话吗?比如朋友难过时该怎么安慰。(Can you teach me a few comforting phrases in Shanghai dialect? For example, what to say when a friend is feeling down.)
Table 4: More Examples Of VCB Bench. Part I.
SIF Size Example
Non Verbal Vocalization 106 说说你鄙视的行为,在说的过程中加入"哼"的鼻音。(Talk about behaviors you despise, and add a "humph" nasal sound while speaking.)
Pacing Control 200 请你慢点读这句 "鱼缸里的水草晃了晃,水珠从叶尖滚下来,慢慢落进水里。"(Please read this sentence slowly: "The aquatic plants in the fish tank swayed slightly, and water droplets rolled off the tips of the leaves, slowly falling into the water.")
Style Control 122 用HR的压迫感提问:"现在,说说你的致命缺点?"(Ask with the pressure of an HR tone: "Now, tell me about your fatal flaw.")
Volume Control 200 聊聊天好吗?用轻柔的语调就行,不用太大声。(Can we chat for a bit? Just use a gentle tone, no need to be too loud.)
@@
Progression 80 Q: 我想开始跑步锻炼,但不知道怎么入门,有什么建议吗?(I want to start running for exercise, but I don’t know how to begin. Any suggestions?)
A: 你可以从慢跑开始,设定小目标,比如每次10分钟,然后逐步增加时间和强度。(You can start with jogging, set small goals, like 10 minutes each time, and then gradually increase the duration and intensity.)
Table 5: More Examples Of VCB Bench. Part II.
MTD Size Example
Progression 80 Q: 为什么要从慢跑开始而不是直接跑得久一些?(Why start with jogging instead of running longer right away?)
A: 慢跑让身体适应运动,减少伤害风险,逐步提高耐力是最有效的方法。(Jogging allows your body to adapt to exercise, reduces the risk of injury, and gradually building endurance is the most effective method.)
Q: 如何知道什么时候该增加跑步时间呢?(How do I know when to increase my running time?)
Backtracking 80 Q: 最近有什么好看的科幻电影推荐吗?(Any good sci-fi movie recommendations lately?)
A: 可以看看《流浪地球》,2023年上映,视觉效果和故事情节都很棒。(You can watch "The Wandering Earth". It was released in 2023 and has great visual effects and storyline.)
Q: 这部电影的故事情节是怎样的?(What is the storyline of this movie about?)
A: 讲述太阳即将毁灭,人类计划移民到别的星球,带着地球一起流浪。(It tells the story of the sun about to be destroyed, with humanity planning to migrate to another planet, taking Earth along on the journey.)
Q: 有没有比较刺激的场面?(Are there any particularly thrilling scenes?)
A: 有地球发动机启动的场景,巨大的动力装置,震撼人心。(There’s the scene where Earth’s engines start up, with massive propulsion devices—it’s awe-inspiring.)
Q: 你刚才说哪一年上映的《流浪地球》?(Which year did you say "The Wandering Earth" was released?)
Transition 80 Q: 最近老是失眠,有什么建议吗?(I’ve been having trouble sleeping lately. Any suggestions?)
A: 晚上试试喝点温牛奶,别吃辛辣食物,放松心情。(Try drinking some warm milk at night, avoid spicy food, and relax.)
Q: 为什么温牛奶有帮助呢?(Why does warm milk help?)
A: 牛奶含色氨酸,有助于睡眠,帮助你放松。(Milk contains tryptophan, which aids sleep and helps you relax.)
Q: 你吃烧烤吗?(Do you eat barbecue?)
@@
Mathematics 36 Q: 二零一八年九月,英国皇家学会前主席迈克尔阿提亚宣称已经解决了哪个世界性的数学难题?(In September 2018, former President of the Royal Society Michael Atiyah claimed to have solved which world-renowned mathematical conjecture?)
A: 黎曼猜想。(The Riemann Hypothesis.)
Geography 150 Q: 我国跨纬度最大的是哪个省级行政区?(Which provincial-level administrative region in China spans the greatest latitude?)
A: 海南省。(Hainan Province.)
Politics 59 Q: 新中国第一部临时宪法的简称是什么?(What is the abbreviated name of New China’s first provisional constitution?)
A: 共同纲领。(The Common Program.)
Table 6: More Examples Of VCB Bench. Part III.